# TIM-PRM: Verifying multimodal reasoning with Tool-Integrated PRM

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have achieved impressive performances in mathematical reasoning, yet they remain vulnerable to visual hallucinations and logical inconsistencies that standard outcome-based supervision fails to mitigate. While Process Reward Models (PRMs) promise step-by-step verification, current approaches typically operate as scalar scorers or generative critics that suffer from sycophancy—blindly validating the flawed hypotheses rather than grounding them in visual reality. To bridge this gap, we introduce TIM-PRM (Tool-Integrated Multimodal PRM), a novel agentic framework that transforms verification from a passive classification task into an active, tool-augmented investigation. TIM-PRM is trained to explicitly plan verification strategies and utilizes a mechanism of Independent Question Asking to query evidence via external tools, effectively decoupling verification from the reasoning context to eliminate confirmation bias. We instantiate this method by curating a high-quality dataset of tool-integrated verification trajectories. Extensive experiments on VisualProcessBench demonstrate that our 8B parameter model surpasses existing open-source multimodal PRMs, significantly outperforming much larger models like Qwen2.5-72B and InternVL-78B, while offering interpretable insights into the verification process.
@@ 1 Introduction
The rapid advancement of Multimodal Large Language Models (MLLMs) has unlocked remarkable capabilities in solving complex reasoning tasks, ranging from geometry [1, 2, 3, 4] and scientific [5] analysis to visual logic puzzles. However, as these models are increasingly deployed in high-stakes domains, ensuring the reliability and faithfulness of their reasoning processes becomes paramount [6, 7]. While standard alignment techniques, such as Reinforcement Learning from Human Feedback (RLHF) [8] with outcome supervision, have improved general response quality, they treat the reasoning chain as a black box. This outcome-focused paradigm often fails to penalize "false positives", where flawed reasoning serendipitously leads to the correct answer, thereby reinforcing hallucinated logic and diminishing the interpretability of the model’s decision-making process.
To address these limitations, recent research has pivoted toward Process Supervision, where a Process Reward Model (PRM) [9, 10, 11] evaluates intermediate steps of a solution. In the multimodal domain, existing approaches typically fall into two categories: discriminative PRMs that output scalar probabilities [7, 12], and generative verifiers that produce linguistic critiques [13, 14]. Despite their promise, both paradigms suffer from certain deficiencies. Scalar PRMs provide "black-box verification," offering no explanation for their scores and struggling to distinguish between subtle visual hallucinations and valid logical leaps. Conversely, generative verifiers often rely solely on the MLLM’s internal parametric knowledge, making them prone to sycophancy: the tendency to blindly agree with the context of the reasoning step rather than critically inspecting the evidence [15]. Consequently, when a reasoning step asserts a false visual fact (e.g., "The graph is a parabola"), a standard verifier is biased to accept this premise, ignoring the error rather than identifying it. This raises the challenge of how to decouple verification from textual reasoning to create a verifier that is not only accurate but also explicitly grounded and resistant to confirmation bias.
@@
  • We identify key failure modes in existing multimodal verification, specifically highlighting the issues of perception hallucinations in current PRMs.
@@ 2 Related works
Multimodal Reward Modeling and Alignment. Reward modeling serves as the cornerstone of aligning Multimodal Large Language Models (MLLMs) with human preferences and ensuring the correctness of generated outputs. Early approaches primarily focused on outcome supervision, where a reward model evaluates the final response quality. Works like InternLM-XComposer2.5-Reward [16] and Skywork-VL Reward [17] pioneered this direction by training discriminative reward models on preference pairs to guide Reinforcement Learning from Human Feedback (RLHF) [8]. However, outcome-based supervision treats the reasoning process as a black box. While effective for general alignment, these methods often fail to locate intermediate logical errors or hallucinations in complex mathematical reasoning tasks, where a correct final answer can occasionally result from flawed reasoning (false positives).
Process Supervision for Multimodal Reasoning. To address the opacity of outcome supervision, recent research has pivoted toward process supervision, which evaluates reasoning trajectories step-by-step. This paradigm has shown significant promise in text-based math reasoning and is now being adapted for multimodal contexts. VisualPRM [7] and URSA [12] introduced discriminative Process Reward Models (PRMs) trained via Monte Carlo Tree Search (MCTS) derived labels. These models assign a scalar probability score to each step, guiding inference-time search algorithms like Best-of-N or Tree Search. Similarly, Athena [18] demonstrated that data-efficient PRMs could be trained by leveraging error-injection techniques. Despite their success, discriminative PRMs suffer from the "black-box verification" problem: they output a score without an explanation. As noted in our analysis (Section 3.1), these models struggle with fine-grained visual grounding errors and often exhibit biases toward simple question replication or outcome-based heuristics. They lack the mechanism to explicitly "look" at the image to verify specific claims.
Generative Verification and Tool Use. Moving beyond scalar scoring, generative verifiers leverage the linguistic capabilities of LLMs to critique reasoning. In the text domain, methods have explored using LLMs to generate natural language feedback or self-corrections. In the multimodal domain, works like MM-RLHF [19], LLaVA-Critic [20], and R1-Reward [21] pioneered in proposing generative reward models. More recently, VRPRM [13] and GM-PRM [14] have begun to explore generative process rewards, enabling PRMs to output reasoning traces alongside verification scores for each reasoning step. However, current generative multimodal verifiers remain constrained by the internal parametric knowledge of the model. They are prone to sycophancy, often agreeing with the solution step’s hallucinated content rather than challenging it. Our work distinguishes itself by integrating explicit tool use into the verification process. Unlike VRPRM or GM-PRM, which rely solely on internal chain-of-thought, TIM-PRM actively queries the visual environment through independent question-asking. This agentic approach decouples perception from reasoning, mitigating the hallucination propagation found in previous generative verifiers.
@@ 3 Revisiting Verification of Multimodal Reasoning
To understand the limitations of current approaches in multimodal mathematical reasoning, we first conduct a preliminary evaluation of existing open-source multimodal PRMs. Through a fine-grained inspection of failure cases, we identify specific challenges related to visual perception and the inherent biases of probability-based reward modeling.
@@ 4.3 Debiasing Sycophantic Verification via Independent Question Asking
$z_{\text{resp}}=\text{Tool}(\mathcal{I},q),\quad\text{where }q\text{ is independent of }h\text{'s truth value}.$ (1)
The verification $v_{t}$ is then derived by comparing the content of $s_{t}$ against the grounded evidence $z_{\text{resp}}$, significantly reducing visual hallucinations. The implementation details and definition of the ask_questions tool are provided in Appendix A.1.
Example Trajectory: We provide an example of how tool is invoked to assist step verification in Figure 4.3.
Tool-Integrated Verification Trajectory ### Paragraph 2 <planning> This paragraph calculates the total sales for Product 1 by summing the values from each store. I need to verify the values from the heatmap and the addition. I will use the ask_questions tool to confirm the values in the heatmap for Product 1 across all stores. </planning> <tool_call> {"name": "ask_questions", "arguments": {"target_image": 1, "questions": ["What are the sales values for Product 1 in Store A, Store B, Store C, and Store D?"]}} </tool_call> <tool> The sales values for Product 1 in Store A, Store B, Store C, and Store D are 119, 177, 116, and 159, respectively. </tool> <analyze> The values for Product 1 are confirmed as 119 (Store A), 177 (Store B), 116 (Store C), and 159 (Store D). The sum is 119 + 177 + 116 + 159 = 571, which matches the calculation in the paragraph. The paragraph is correct. </analyze> <verify> CORRECT </verify>
@@ 5 Experiments
5.1 Experiment Setup
Benchmarks.
To evaluate the efficacy of TIM-PRM in verifying multimodal mathematical reasoning, we employ VisualProcessbench as the main benchmark, which contains reasoning trajectories from a diverse set of policy models on 5 benchmarks with human-annotated step-level correctness. Specifically, we report the subset results on MMMU [5], MathVision [1], MathVerse-VO [4], DynaMath [3], and WeMath [2]. These datasets cover a wide spectrum of visual reasoning tasks, ranging from geometry and function analysis to complex multi-step logical deduction. We specifically focus on the subsets of these benchmarks that require step-by-step reasoning, adhering to the standard evaluation protocols defined in previous process supervision literature. We consider the neutral steps as a correct step following Lightman et al. [6].
Baselines.
  • Proprietary MLLMs: We evaluate state-of-the-art closed-source models including GPT-4o [22], GPT-4o-Mini [22], and Gemini-2.0-Flash [23]. These models are prompted to act as verifiers (MLLM-as-a-PRM) using zero-shot instructions to critique the correctness of reasoning steps [7].
  • Open-source MLLMs: We include leading open-weights models such as MiniCPM-V2.6 [24], LLaVA-OV [25], Qwen2.5-VL [26], InternVL2.5 [27], and Qwen3-VL [28] across various parameter scales (ranging from 7B to 78B). Similar to the proprietary models, these are evaluated in a generative verifier setting.
  • Specialized Multimodal PRMs: We compare against VisualPRM-8B [7] and $M$-PRM-8B [29], which are trained specifically for process reward modeling using traditional outcome-based supervision or standard preference optimization techniques without tool integration.
Training data.
We use the MathV360K subset of the VisualPRM400K dataset because it ensembles a variety of datasets, containing about 20.1k samples. After format filtering and consensus filtering [30] with MLLM-as-a-PRM implemented by Qwen3-VL-30B-A3B-Instruct, about 13k samples with reliable process labels remains, with a pass rate of 65%.
Implementation Details.
5.2 Main results
Step-wise Verification Performance.
Table 2: Performance comparison of various models on macro F1 score. The results of models marked with are cited from [7]
@@ 5.2 Main results
TIM-PRM-2B 55.8 57.6 61.7 64.3 58.0 60.3
TIM-PRM-8B 58.3 58.3 61.9 65.9 63.9 61.7
First Incorrect Step Identification.
@@ 5.3 Ablation Studies
Figure 1: Analysis of Data Filtering Consistency.
@@ A.1 Tool Definitions
The prompt in Figure A.2 defines the ‘ask_questions‘ tool used by the model. It is designed to query an external MLLM to gather visual information without the bias of the solution text.
@@ A.2 Prompt Structure for Trajectory Generation
We define the prompt for generating tool-calling trajectories in Figure A.2. This structure ensures the model adheres to the <planning> $\to$ <tool_call> $\to$ <analyze> $\to$ <verify> workflow.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have achieved impressive performances in mathematical reasoning, yet they remain vulnerable to visual hallucinations and logical inconsistencies that standard outcome-based supervision fails to mitigate. While Process Reward Models (PRMs) promise step-by-step verification, current approaches typically operate as scalar scorers or generative critics that suffer from sycophancy, blindly validating the flawed hypotheses rather than grounding them in visual reality. To bridge this gap, we introduce TIM-PRM (Tool-Integrated Multimodal PRM), a novel agentic framework that transforms verification from a passive classification task into an active, tool-augmented investigation. TIM-PRM is trained to explicitly plan verification strategies and utilizes a mechanism of Independent Question Asking to query evidence via external tools, effectively decoupling verification from the reasoning context to eliminate confirmation bias. We instantiate this method by curating a high-quality dataset of tool-integrated verification trajectories. Extensive experiments on VisualProcessBench demonstrate that our 8B parameter model surpasses existing open-source multimodal PRMs, significantly outperforming much larger models like Qwen2.5-72B and InternVL-78B, while offering interpretable insights into the verification process.
TIM-PRM: Verifying multimodal reasoning with Tool-Integrated PRM
Peng Kuang1,2, Xiangxiang Wang2, Wenchao Liu2, Jian Dong2, Kaidi Xu3 1Zhejiang University, 2iFLYTEK AI Research Institute, 3City University of Hong Kong, pengkuang@zju.edu.cn, {xxwang56,wcliu4,jiandong3}@iflytek.com, kaidixu@cityu.edu.hk
@@ 1 Introduction
The rapid advancement of Multimodal Large Language Models (MLLMs) has unlocked remarkable capabilities in solving complex reasoning tasks, ranging from geometry Wang et al. (2024); Qiao et al. (2025); Zou et al. (2024); Zhang et al. (2024) and scientific Yue et al. (2024) analysis to visual logic puzzles. However, as these models are increasingly deployed in high-stakes domains, ensuring the reliability and faithfulness of their reasoning processes becomes paramount Lightman et al. (2023); Wang et al. (2025b). While standard alignment techniques, such as Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022) with outcome supervision, have improved general response quality, they treat the reasoning chain as a black box. This outcome-focused paradigm often fails to penalize "false positives", where flawed reasoning serendipitously leads to the correct answer, thereby reinforcing hallucinated logic and diminishing the interpretability of the model’s decision-making process.
To address these limitations, recent research has pivoted toward Process Supervision, where a Process Reward Model (PRM) (Lightman et al., 2024; Li and Li, 2025; Zheng et al., 2024) evaluates intermediate steps of a solution. In the multimodal domain, existing approaches typically fall into two categories: discriminative PRMs that output scalar probabilities Wang et al. (2025b); Luo et al. (2025), and generative verifiers that produce linguistic critiques noa (2025); Zhang et al. (2025a). Despite their promise, both paradigms suffer from certain deficiencies. Scalar PRMs provide "black-box verification," offering no explanation for their scores and struggling to distinguish between subtle visual hallucinations and valid logical leaps. Conversely, generative verifiers often rely solely on the MLLM’s internal parametric knowledge, making them prone to sycophancy: the tendency to blindly agree with the context of the reasoning step rather than critically inspecting the evidence Liu et al. (2025). Consequently, when a reasoning step asserts a false visual fact (e.g., "The graph is a parabola"), a standard verifier is biased to accept this premise, ignoring the error rather than identifying it. This raises the challenge of how to decouple verification from textual reasoning to create a verifier that is not only accurate but also explicitly grounded and resistant to confirmation bias.
Figure 1: Comparison between traditional passive verification and TIM-PRM. (a) Existing PRMs act as passive scorers, suffering from sycophancy where they blindly validate hallucinated claims due to reliance on text context. (b) TIM-PRM transforms verification into an agentic task, employing explicit planning, independent tool querying to decouple perception from reasoning, and grounded analysis to detect visual hallucinations.
@@
  • We identify key failure modes in existing multimodal verification, highlighting the issues of perception hallucinations in current PRMs.
@@ 2 Related works
Process Supervision for Multimodal Reasoning. To address the opacity of outcome supervision, recent research has pivoted toward process supervision, which evaluates reasoning trajectories step-by-step. This paradigm has shown significant promise in text-based math reasoning and is now being adapted for multimodal contexts. VisualPRM Wang et al. (2025b) and URSA Luo et al. (2025) introduced discriminative Process Reward Models (PRMs) trained via Monte Carlo Tree Search (MCTS) derived labels. These models assign a scalar probability score to each step, guiding inference-time search algorithms like Best-of-N or Tree Search. Similarly, Athena Wang et al. (2025a) demonstrated that data-efficient PRMs could be trained by leveraging error-injection techniques. Despite their success, discriminative PRMs suffer from the "black-box verification" problem: they output a score without an explanation. As noted in our analysis (Section 3.1), these models struggle with fine-grained visual grounding errors and often exhibit biases toward simple question replication or outcome-based heuristics. They lack the mechanism to explicitly "look" at the image to verify specific claims.
Generative Verification and Tool Use. Moving beyond scalar scoring, generative verifiers leverage the linguistic capabilities of LLMs to critique reasoning. In the text domain, methods have explored using LLMs to generate natural language feedback or self-corrections. In the multimodal domain, works like MM-RLHF Zhang et al. (2025c), LLaVA-Critic Xiong et al. (2025), and R1-Reward Zhang et al. (2025b) pioneered in proposing generative reward models. More recently, VRPRM noa (2025) and GM-PRM Zhang et al. (2025a) have begun to explore generative process rewards, enabling PRMs to output reasoning traces alongside verification scores for each reasoning step. However, current generative multimodal verifiers remain constrained by the internal parametric knowledge of the model. They are prone to sycophancy, often agreeing with the solution step’s hallucinated content rather than challenging it. Our work distinguishes itself by integrating explicit tool use into the verification process. Unlike VRPRM or GM-PRM, which rely solely on internal chain-of-thought, TIM-PRM actively queries the visual environment through independent question-asking. This agentic approach decouples perception from reasoning, mitigating the hallucination propagation found in previous generative verifiers.
@@ 3 Revisiting Verification of Multimodal Reasoning
To understand the limitations of current approaches in verifing multimodal reasoning, we first conduct a preliminary evaluation of existing open-source multimodal PRMs. Through a fine-grained inspection of failure cases, we identify specific challenges related to visual perception and the inherent biases of probability-based reward modeling.
@@ 4.3 Debiasing Sycophantic Verification via Independent Question Asking
$z_{\text{resp}}=\text{Tool}(\mathcal{I},q)$ (1)
where $q$ is independent of $h$’s truth value. The verification $v_{t}$ is then derived by comparing the content of $s_{t}$ against the grounded evidence $z_{\text{resp}}$, significantly reducing visual hallucinations. The implementation details of the ask_questions tool are provided in Appendix A.1.
Example Trajectory: We provide an example of tool invoking in step verification in Figure A.1.
@@ 5 Experiments
Table 2: Performance comparison of various models on macro F1 score. The results of models marked with are cited from Wang et al. (2025b)
@@
[…]
5.1 Experiment Setup
Benchmarks.
To evaluate the efficacy of TIM-PRM in verifying multimodal mathematical reasoning, we employ VisualProcessbench as the main benchmark, which contains reasoning trajectories from a diverse set of policy models on 5 benchmarks with human-annotated step-level correctness. Specifically, we report the subset results on MMMU Yue et al. (2024), MathVision Wang et al. (2024), MathVerse-VO Zhang et al. (2024), DynaMath Zou et al. (2024), and WeMath Qiao et al. (2025). These datasets cover a wide spectrum of visual reasoning tasks, ranging from geometry and function analysis to complex multi-step logical deduction. We specifically focus on the subsets of these benchmarks that require step-by-step reasoning, adhering to the standard evaluation protocols defined in previous process supervision literature. We consider the neutral steps as a correct step following Lightman et al. (2023).
Baselines.
  • Proprietary MLLMs: We evaluate state-of-the-art closed-source models including GPT-4o noa , GPT-4o-Mini noa , and Gemini-2.0-Flash Team (2024). These models are prompted to act as verifiers (MLLM-as-a-PRM) using zero-shot instructions to critique the correctness of reasoning steps Wang et al. (2025b).
  • Open-source MLLMs: We include leading open-weights models such as MiniCPM-V2.6 Yao et al. (2024), LLaVA-OV Li et al. (2024), Qwen2.5-VL Bai et al. (2025b), InternVL2.5 Chen et al. (2025), and Qwen3-VL Bai et al. (2025a) across various sizes ranging from 7B to 78B. Similar to the proprietary models, these are evaluated in a generative verifier setting.
  • Specialized Multimodal PRMs: We compare against VisualPRM-8B Wang et al. (2025b) and $M$-PRM-8B Du et al. (2025), which are trained specifically for process reward modeling using traditional outcome-based supervision or standard preference optimization techniques without tool integration.
Training data.
We use the MathV360K subset of the VisualPRM400K dataset because it ensembles a variety of datasets, containing about 20.1k samples. After format filtering and consensus filtering Zhang et al. (2025d) with MLLM-as-a-PRM implemented by Qwen3-VL-30B-A3B-Instruct, about 13k samples with reliable process labels remains, with a pass rate of 65%.
Implementation Details.
@@ 5.1 Experiment Setup
5.2 Main results
Step-wise Verification Performance.
First Incorrect Step Identification.
5.3 Ablation Studies
@@
(a) Analysis of First Error Identification.
(b) Analysis of Data Filtering Consistency.
Figure 2: Training data analysis.
@@ 6 Conclusion
Limitations
While TIM-PRM demonstrates significant improvements in verifying multimodal reasoning, there are avenues for further optimization. First, as an agentic framework involving explicit planning and tool interaction, our approach inherently incurs a higher inference cost compared to passive scalar verifiers. Future work could address this by exploring knowledge distillation techniques to transfer the verification capabilities into a smaller model, or by implementing adaptive verification strategies that only trigger the agentic workflow for highly uncertain steps. Second, our current implementation focuses primarily on visual verification via Independent Question Asking. While effective for perception, this does not fully address complex logical or calculation errors; integrating code interpreters or symbolic solvers could further enhance logical rigor. Finally, while we validate our method on diverse mathematical benchmarks, exploring its application to broader multimodal domains, such as general visual scene understanding or document analysis, remains a critical direction for future research.
@@ A.1 Tool Definitions
The prompt in Figure 3 defines the ‘ask_questions‘ tool used by the model. It is designed to query an external MLLM to gather visual information without the bias of the solution text.
Verification Trajectory with Tools ### Paragraph 2 <planning> This paragraph calculates the total sales for Product 1 by summing the values from each store. I need to verify the values from the heatmap and the addition. I will use the ask_questions tool to confirm the values in the heatmap for Product 1 across all stores. </planning> <tool_call> {"name": "ask_questions", "arguments": {"target_image": 1, "questions": ["What are the sales values for Product 1 in Store A, Store B, Store C, and Store D?"]}} </tool_call> <tool> The sales values for Product 1 in Store A, Store B, Store C, and Store D are 119, 177, 116, and 159, respectively. </tool> <analyze> The values for Product 1 are confirmed as 119 (Store A), 177 (Store B), 116 (Store C), and 159 (Store D). The sum is 119 + 177 + 116 + 159 = 571, which matches the calculation in the paragraph. The paragraph is correct. </analyze> <verify> CORRECT </verify>
@@ A.2 Prompt Structure for Trajectory Generation
We define the prompt for generating tool-calling trajectories in Figure 3. This structure ensures the model adheres to the <planning> $\to$ <tool_call> $\to$ <analyze> $\to$ <verify> workflow.